the height of a tower is 3 times the height of a certain building. if the tower is 50m taller than the building,how tall is the tower?
Answer by checkley77(12844)

Tower=3*building
Tower=building+50
3b=b+50
3b-b=50
2b=50
b=50/2
b=25 m tall.
Proof:
t=3*25=75m for the tower.
t=25+50=75 for the tower.